# Large Language Models:

### v1

@@ 3.1 Setting and the Researcher’s Dataset
Let $\Sigma^{*}$ denote the collection of strings (up to some finite length) in an alphabet with elements $\sigma\in\Sigma^{*}$, and a training dataset is any collection of strings. We summarize a training dataset by the vector $t$, whose elements $t_{\sigma}$ are sampling indicators for whether a particular string is collected in the training dataset.
Of course, for any given empirical question, not all strings are relevant; we denote those that are as $\mathcal{R}\subseteq\Sigma^{*}$ with elements $r\in\mathcal{R}$ that we refer to as text pieces. The researcher’s dataset is also summarized by the vector $d$, whose elements $d_{\sigma}$ are similarly sampling indicators for whether the researcher collected a particular string. The researcher only collects economically relevant text pieces, and so $d_{\sigma}=0$ for all strings $\sigma\in\Sigma^{*}\setminus\mathcal{R}$.
Each text piece $r$ is (or can be) linked to observable economic variables $(Y_{r},W_{r})$, which can be thought of as economic outcomes that might be influenced by the text ($Y_{r}$) or candidate economic determinants that might influence the text ($W_{r}$). Altogether the researcher observes $(r,Y_{r},W_{r})$ for each collected text piece with $d_{r}=1$. To make this more concrete, consider two illustrative empirical applications that we return to throughout the paper.
@@
Consider descriptions of bills introduced in the United States Congress. Each text piece $r\in\mathcal{R}$ refers to a bill’s brief description such as “A bill to revise the boundary of Crater Lake National Park in the State of Oregon.” The economically relevant outcome $Y_{r}$ might be whether the associated bill passed its originating chamber of Congress. The candidate economic determinant $W_{r}$ of the bill’s text might be the party affiliation or roll-call voting score of the bill’s sponsor. $\blacktriangle$
@@
Consider financial news headlines about publicly traded companies. Each text piece $r\in\mathcal{R}$ refers to a particular financial news headline such as “Bank of New York Mellon Q1 EPS \$0.94 Misses \$0.97 Estimate, Sales \$3.9B Misses \$4.01B Estimate.” The economic outcome $Y_{r}$ might be the company’s realized return in some event window after the headline’s publication date, while the candidate determinant $W_{r}$ of the news headline itself could be the company’s past fundamentals. $\blacktriangle$
Importantly, each text piece $r$ also expresses some economic concept $V_{r}$ for which it is costly in terms of either time or money to obtain measurements. A key assumption (often left implicit) in much empirical research is that there is some procedure that could in principle be applied to each text piece to collect what the researcher would be willing to consider a “gold standard” measurement of the economic concept; that is, there exists some mapping such that $V_{r}=f^{*}(r)$ for all text pieces $r\in\mathcal{R}$.
For example, in some applications a researcher is willing to argue that a sufficiently reliable measure of the economic concept could be derived by having a single expert spend the time and effort needed to read each text piece carefully. [^6] For example, AshHansen(23) write, “The most accurate approach to concept detection is perhaps direct human reading with appropriate domain expertise” (pg. 672). HansenEtAl(23)-RemoteWork write, “The most precise way of classifying [text pieces] is arguably via direct human reading” (pg. 6). In other applications, the researcher, motivated by findings in psychology (e.g., biemer2013measurement; kahneman2021noise), will worry that human annotations of some text are noisy proxies for the economic concept, but is willing to accept an average from some minimum number of labelers as a reliable measure.
@@
No matter how the status quo procedure for collecting measurements is defined, we face a text processing problem: measuring the economic concept $V_{r}$ requires processing each text piece $r$, and it may be prohibitively costly or time-consuming to do so in many applications. Absent a solution to this text processing problem, the researcher does not observe $V_{r}$.
In settings like this, researchers would like to process the collected text pieces to tackle one of two types of economic analyses. The first is a prediction problem — predict the linked variable $Y_{r}$ using the associated text piece $r$. For example, we might use the short text description of some Congressional legislation $r$ to predict whether the bill passed its originating chamber of Congress $Y_{r}$. Or a researcher might try to use each financial news headline $r$ to predict the company’s realized return $Y_{r}$ in some event window after the headline’s publication date.
The second is an estimation problem — estimate some parameter that relates the economic concept $V_{r}$ to the linked variables $(Y_{r},W_{r})$. For example, if each Congressional bill’s brief text description $r$ expresses the policy topic $V_{r}$ of the bill, such as whether it is related to defense, foreign affairs, or health, that topic could be related in a regression to the party affiliation or ideological voting score of the bill’s sponsor $W_{r}$. Or if each financial news headline $r$ expresses positive or negative news about some company’s future $V_{r}$, that could be the outcome in a regression against past fundamentals $W_{r}$ or the explanatory variable in a regression against future returns.
@@
This is where LLMs come in. LLMs are general-purpose and easy-to-use models for processing text, so researchers would like to use them to tackle their text processing challenge in service of some prediction or estimation problem. When is that justified? To clarify those conditions, we next introduce LLMs into our framework.
3.2 Large Language Models
@@
To capture how we typically interact with LLMs as black boxes — to generate responses given prompts without knowing or worrying much about their design or the contents of the training datasets — we define a large language model as any mapping from possible training datasets $t$ to mappings between strings, where $\widehat{m}(\cdot;t)\colon\Sigma^{*}\rightarrow\Sigma^{*}$ is its text generator when it is trained on dataset $t$ and $\widehat{m}(\sigma;t)$ is the LLM’s response when prompted by string $\sigma$.
@@
While this definition is quite general, notice it has a specific implication: The LLM “algorithm” is actually two algorithms: a training algorithm that takes in any training dataset $t$ and learns the mapping between strings; and what we call the text generator, $\widehat{m}(\cdot;t)$, which is the output of the training algorithm and is what the user interacts with to obtain responses to any given prompt. (We return below to the importance of this implication).
The other important feature of our definition is that our analysis will not depend on how exactly an LLM’s training algorithm and text generator accomplish their functionalities. This is intentional. We would like to understand the conditions under which we can use any given LLM to solve prediction and estimation problems. By analyzing them at this level of abstraction, we provide interpretable conditions on the LLM needed for empirical researchers to accomplish their objectives.
@@
Furthermore, since the state of the art in natural language processing is constantly evolving, we should expect the exact implementation of training algorithms and text generators to change; studying LLMs at this level of abstraction ensures the durability of our analysis. In our framework, alternative LLMs, such as GPT-3.5-Turbo and GPT-4o versus Llama-3-8B and Llama-3-70B (or alternative snapshots of the same model), may all differ in their training algorithms and text generators, but researchers will now have a clear set of consistent conditions under which the output of any given model can (or cannot) be incorporated into an economics research pipeline.
3.2.1 Interpreting the Text Generator and the Training Algorithm
Our framework, even as general as it is, nonetheless captures all of the key design choices underlying LLMs. Some of those choices are made by the researcher themselves, while others are made by the algorithm builder — many of which may even be invisible to the researcher.
@@ 3.2.1 Interpreting the Text Generator and the Training Algorithm
The researcher interacts entirely with the text generator $\widehat{m}(\cdot;t)$, which winds up capturing all choices that influence how an LLM generates responses from any prompt once it has been trained. For example, research in computer science has found that alternative prompt engineering strategies, such as personas, chain-of-thought reasoning, or few-shot prompting, materially affect the quality of responses given by an LLM to a fixed task or question (LiuEtAl(23); wei2023chainofthoughtpromptingelicitsreasoning; white2023promptpatterncatalogenhance; chen2024unleashingpotentialpromptengineering). In our framework, any particular prompt engineering strategy can be cast as an alternative choice of the text generator $\widehat{m}(\cdot;t)$.
The text generator additionally captures alternative choices of other hyper-parameters that govern how the LLM randomly generates text in response to a given prompt, such as its “temperature,” top-$p$ sampling or top-$k$ sampling parameters. An LLM models the probability distribution over tokens that are likely to come next given an existing string; these hyperparameters all govern the extent of randomness in how the text generator samples from its probability distribution over next tokens. Alternative LLMs may have different default choices of these parameters; various APIs give researchers access to choose these parameters themselves. By defining the text generator as a deterministic mapping from prompts to responses, our framework can be interpreted as focusing on the case in which these parameters are such that the LLM greedily only generates its most likely token. Our results extend naturally to stochastic text generators as well, but at the expense of more cumbersome notation.
We briefly note that the text generator $\widehat{m}(\cdot;t)$ also captures other design choices that are typically hidden to the user. Since it is computationally costly to generate text from the transformer architecture, there exist alternative “inference algorithms” that may be used, such as greedy decoding or beam search. Since alternative inference algorithms may affect an LLM’s outputs, alternative choices correspond to alternative text generators in our framework.
Finally, the training algorithm captures all aspects of the design and training process of an LLM that produces the text generator. In particular, it captures the architecture, such as the total number of parameters, the number of attention layers, its choice of proxy objective in pre-training (such as next token prediction or variants like masked token prediction), its context window, and the details of its optimization procedure (batch size, initialization strategies, etc.). We emphasize that the training algorithm captures both pre-training and any additional fine-tuning, such as instruction tuning and reinforcement learning from human feedback. In this sense, the training dataset $t$ in our framework must be interpreted as containing all strings upon which the LLM was pre-trained and fine-tuned on.
4 Prediction with LLMs
With this framework in hand, we can turn to identifying the conditions under which the LLM can be incorporated into an economics research pipeline. We start with what we defined above as a “prediction problem.” A researcher would like to assess the predictability of the linked economic variable $Y_{r}$ based on the text pieces $r$. The economic variable could be, say, a stock price or an indicator for whether some piece of legislation passed Congress, but (as we discuss further below) could also be something more abstract like a new scientific hypothesis.
While the researcher could in principle build their own predictor based on text from scratch, doing so would require vast amounts of training data to first learn the sort of lower-dimensional representation of language needed to successfully predict from text. Since LLMs have already learned some lower-dimensional representation, having been trained on enormous datasets, we may hope to use them as a foundation upon which to tackle our prediction problem.
We use our framework to establish the key condition under which the use of an LLM for this type of research question is valid: there must be no leakage between the LLM’s training dataset and the researcher’s context. Training leakage is a thorny problem for benchmark evaluations of LLMs in computer science, and it places meaningful limits on the types of prediction problems that researchers can tackle using LLMs. Nonetheless, we discuss how training leakage can be managed through careful choice of the researcher’s context and LLM. As a result, LLMs can be valuable tools for solving prediction problems, provided we take the threat of training leakage seriously — far more seriously than has typically been the case in social science research to date.
@@ 4.1 The Researcher’s Prediction Problem
Suppose the researcher prompts an LLM, trained on some unknown training dataset $t$, to form predictions of the linked variable based on the text pieces, $\widehat{Y}_{r}=\widehat{m}(r;t)$. [^7] Researchers may instead use a given LLM to construct embeddings for each text piece $r$, and the resulting embeddings may then be used as features by a supervised machine learning algorithm to predict $Y_{r}$. Our analysis in this section equally applies to evaluating the performance of a prediction function that uses LLM embeddings as features. For some non-negative loss function $\ell(y,\tilde{y})$, the researcher then calculates the sample average loss of the LLM’s predictions on their collected dataset:
@@
where $N=\sum_{r}D_{r}$ is the number of text pieces collected by the researcher. We would like to draw conclusions about the predictability of the linked economic variable $Y_{r}$ from the text piece $r$ based on the LLM’s sample average loss (Equation 1). What conditions must the LLM satisfy in order for this to be valid?
To answer this question, we define the researcher’s inferential goal in the prediction problem. We associate the researcher with a context $Q(\cdot)\in\mathcal{Q}$ that summarizes their own sampling distribution over economically relevant text pieces and beliefs about the LLM’s sampling distribution over training datasets; more precisely, $Q(\cdot)$ is a joint distribution over the sampling indicators $(D,T)$. The researcher’s context $Q(\cdot)$ therefore summarizes two distinct features: first, it defines the collection of text pieces over which the researcher would like to assess the predictability of $Y_{r}$ — this is chosen and known to the researcher; second, since its exact contents are unknown, it also captures the researcher’s uncertainty over what strings entered into the LLM’s training dataset.
We assume that the collection of research contexts $\mathcal{Q}$ satisfies the following assumptions.
@@
Letting $t=\left(t_{\sigma_{1}},\ldots,t_{\sigma_{|\Sigma^{*}|}}\right)^{\prime}$ denote the large language model’s realized training dataset, all research contexts $Q(\cdot)\in\mathcal{Q}$ satisfy the following assumptions:
  1. (i)
    For all values $d$, $Q(D=d,T=t)=\Pi_{\sigma\in\Sigma^{*}}Q(D_{\sigma}=d_{\sigma},T_{\sigma}=t_{\sigma})$.
  2. (ii)
    $\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}]=\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\mid T=t]$.
Assumption 1(i) states that the researcher’s and the LLM’s sampling process over strings is independent but not identically distributed, capturing the idea that not all strings may be sampled with equal probability. Assumption 1(ii) states that irrespective of the LLM’s corpus, the researcher always samples the same number of text pieces on average. As notation, $q_{\sigma}^{T\mid D}(t_{\sigma})=Q(T_{\sigma}=t_{\sigma}\mid D_{\sigma}=1)$ is the conditional probability the string is sampled by the LLM’s training dataset given that it is sampled by the researcher. We let $q_{\sigma}^{T}(t_{\sigma})=Q(T_{\sigma}=t_{\sigma})$ and $q_{\sigma}^{D}=Q(D_{\sigma}=1)$.
Conditional on the LLM’s realized training dataset, under mild regularity conditions as the number of economically relevant text pieces grows large (see Appendix C.1), the researcher’s sample average loss converges to

### v3

@@ 2.1 Setting and the Researcher’s Dataset
Let $\Sigma^{*}$ denote a collection of strings (up to some finite length) in an alphabet with elements $\sigma\in\Sigma^{*}$. A training dataset is any collection of strings, summarized by the vector $t$ whose elements $t_{\sigma}$ are sampling indicators for whether string $\sigma$ was collected.
For any empirical question, only some strings are economically relevant. Denote these as $\mathcal{R}\subseteq\Sigma^{*}$ with elements $r\in\mathcal{R}$ that we refer to as text pieces. The researcher’s dataset is summarized by the vector $d$ whose elements $d_{\sigma}$ are sampling indicators for whether the researcher collected string $\sigma$. The researcher only collects economically relevant text pieces, and so $d_{\sigma}=0$ for all $\sigma\in\Sigma^{*}\setminus\mathcal{R}$.
Each text piece $r$ can be linked to observable economic variables $(Y_{r},W_{r})$, which are economic outcomes $Y_{r}$ that might be influenced by the text or candidate covariates $W_{r}$ that might influence the text.
@@
Consider descriptions of bills introduced in the United States Congress. Each text piece $r\in\mathcal{R}$ is a bill’s description such as “A bill to revise the boundary of Crater Lake National Park in the State of Oregon.” The economic outcome $Y_{r}$ might be whether the bill passed its originating chamber. The covariate $W_{r}$ might be the party affiliation or roll-call voting score of the bill’s sponsor. $\blacktriangle$
@@
Consider financial news headlines about publicly traded companies. Each text piece $r\in\mathcal{R}$ is a financial news headline such as “Bank of New York Mellon Q1 EPS \$0.94 Misses \$0.97 Estimate, Sales \$3.9B Misses \$4.01B Estimate.” The economic outcome $Y_{r}$ might be the company’s realized return in some window after the headline’s publication date, while the covariate $W_{r}$ could be the company’s past fundamentals. $\blacktriangle$
Each text piece $r$ expresses some economic concept $V_{r}:=f^{*}(r)$, where $f^{*}(\cdot)$ is the measurement procedure the researcher would use if time and resources were no constraint. Typically these measurements are what the researcher themselves or an appropriate domain expert would produce by carefully reading each text piece. [^2] For example, AshHansen(23) write, “The most accurate approach to concept detection is perhaps direct human reading with appropriate domain expertise” (p. 672). HansenEtAl(23)-RemoteWork write, “The most precise way of classifying [text pieces] is arguably via direct human reading” (p. 6). In defining the economic concept, researchers must answer: what exactly am I measuring from the text, and how would I label it if resources were no constraint? Moving forward, we assume a measurement procedure $f^{*}(\cdot)$ exists that the researcher would be satisfied to use if it could be scaled.
@@
This creates a text processing problem: measuring the economic concept requires processing each text piece $r$, which may be prohibitively costly at scale. Absent a solution to this text processing problem, the researcher cannot collect $V_{r}$ on all text pieces.
In settings like this, researchers would like use the collected text pieces to tackle two types of empirical analyses. The first is a prediction problem: predict the linked variable $Y_{r}$ using the associated text piece $r$. For example, can we predict stock returns from news headlines? The second is an estimation problem: estimate some parameter that relates the economic concept $V_{r}$ to the linked variables $(Y_{r},W_{r})$. If bill descriptions $r$ express the policy topic $V_{r}$ of the bill (e.g., whether it is related to foreign affairs or health policy), how do policy topics relate to the party affiliation of the bill’s sponsor $W_{r}$?
@@
LLMs are general-purpose and easy-to-use models for processing text, so researchers would like to use them in prediction and estimation problems. We next introduce LLMs into our framework.
2.2 Incorporating Large Language Models in Empirical Research
@@
To capture how we often interact with LLMs as black boxes — generating responses from prompts without knowing their design nor training data — we define a large language model as any mapping from possible training datasets $t$ to mappings between strings, where $\widehat{m}(\cdot;t)\colon\Sigma^{*}\rightarrow\Sigma^{*}$ is its text generator when trained on dataset $t$ and $\widehat{m}(\sigma;t)$ is the LLM’s response when prompted by string $\sigma$.
@@
This definition has a key implication: the LLM is actually two distinct algorithms. First, a training algorithm that takes any dataset $t$ and learns the mapping between strings. Second, the text generator $\widehat{m}(\cdot;t)$ is the output of the training algorithm and what users interact with.
@@
Our analysis will not depend on how exactly these algorithms work. Since the state of the art is constantly evolving, we should expect the exact implementation of training algorithms and text generators to change. Furthermore, alternative LLMs may differ in their implementations. Studying LLMs at this level of abstraction will provide interpretable conditions for empirical research and ensures the durability of our analysis. Researchers will have conditions under which any model’s output can be incorporated into empirical research.
Our framework captures all key LLM design choices — some made by researchers, others by algorithm builders (often invisibly to researchers).
Interpreting the Text Generator
@@
The text generator $\widehat{m}(\cdot;t)$ captures all choices that influence how an LLM generates responses once trained. This includes prompt engineering strategies that materially affect the quality of responses (e.g., LiuEtAl(23); wei2023chainofthoughtpromptingelicitsreasoning; white2023promptpatterncatalogenhance; chen2024unleashingpotentialpromptengineering). Alternative prompt engineering strategies can be cast as alternative specifications of the text generator $\widehat{m}(\cdot;t)$.
The text generator captures other important choices governing generation. Parameters like temperature, top-$p$ sampling, or top-$k$ sampling control randomness in sampling from the LLM’s probability distribution over next tokens. Reasoning models employ test-time computation, generating intermediate “thought tokens” before producing final outputs to enhance performance on complex tasks. [^3] For many reasoning models, users cannot control generation parameters like temperature. Consequently, there is inherent randomness that cannot be eliminated. By defining the text generator as a deterministic mapping from prompts to responses, our framework can be interpreted as focusing on the case in which these parameters are such that the LLM greedily generates its most likely token. Our results extend naturally to stochastic text generators as well, but at the expense of more cumbersome notation. [^4] A practical challenge is ensuring LLM outputs are formatted appropriately for empirical analysis. Researchers often prompt LLMs to return structured outputs, but LLMs may not reliably comply. Our text generator abstraction takes further post-processing steps as given, and $\widehat{m}(\cdot;t)$ represents the resulting final output.
Interpreting the Training Algorithm
The training algorithm captures all aspects of design and training that produce the text generator. This includes architectural choices (e.g., parameter count, attention layers, context window), the pretraining objective (typically next-token prediction), and optimization details. LLMs undergo multiple training stages beyond pretraining. Instruction fine-tuning trains models to follow user instructions, and reinforcement learning from human feedback (RLHF) aligns output with human preferences. More recent developments include reinforcement learning from verifiable rewards (RLVR), where models are trained on tasks with objectively correct answers. Crucially, the training dataset $t$ in our framework encompasses all strings used in pretraining and post-training stages.
So how do researchers use LLMs in empirical research? In prediction problems, researchers often prompt an LLM with each text piece $r$ to generate a prediction $\hat{Y}_{r}=\widehat{m}(r;t)$. For instance, they might prompt a model to predict whether a stock return will be positive. In estimation problems, researchers prompt an LLM to generate labels $\hat{V}_{r}=\widehat{m}(r;t)$ for the economic concept. For example, they might prompt a model to classify a bill’s policy topic.
2.3 The Challenge of Evaluating Large Language Models
At some level, the viability of using LLMs for prediction and estimation problems hinges on the quality of LLM outputs: how large are possible errors $Y_{r}-\widehat{Y}_{r}$ and $V_{r}-\widehat{V}_{r}$ in any application? A natural starting point is to understand how computer science approaches this evaluation problem.
Computer scientists adapted methods from supervised learning. In supervised learning, the “common task framework” (Donoho2024Data) builds benchmark datasets like ImageNet for image classification and the Netflix Prize for recommender systems and evaluates competing algorithms on these benchmarks. Since LLMs aspire to be general-purpose technologies useful across all tasks, computer scientists have built increasingly diverse benchmarks. For example, the ”Beyond-the-Imitation-Game benchmark” (BIG-bench) collects problems across 204 tasks (srivastava2022beyond), while the ”Massive Multitask Language Understanding” (MMLU) benchmark spans 57 categories from logic to social sciences (hendrycks2020measuring). Other benchmarks assess specialized capabilities: SWE-bench for coding ability (jimenez2024swebenchlanguagemodelsresolve), GSM8K for mathematical reasoning (cobbe2021trainingverifierssolvemath), and standardized exams such as the SAT, GRE, and AP tests. Modern LLMs perform remarkably well on these benchmarks, and this impressive performance forms much of the quantitative basis for current enthusiasm.
Can economists use these benchmark evaluations to reason about how LLMs will perform on specific empirical tasks? The answer is more complicated than it might initially appear.
2.3.1 The Limits of Benchmarks
We do not intrinsically care about benchmark performance itself — after all, who deploys LLMs to solve SAT problems? We hope to generalize from benchmarks to new economically relevant tasks.
This generalization happens intuitively. We assume an LLM that aces AP chemistry must, like a person, handle many chemistry-related tasks. Evidence for “anthropomorphic generalization” suggests that people apply similar generalization heuristics to LLMs as they do to humans when predicting performance across tasks (vafa2024humangeneralization; DreyfussRaux(24)).
Yet this intuition is misleading. LLMs exhibit what mancoridis2025potemkin call “potemkin understanding” — performing well on benchmarks without grasping underlying concepts. This manifests as remarkable brittleness: performance is sensitive to seemingly minor details that would not affect human performance. Consider several examples from an accumulating body of evidence. While humans who master one math problem typically solve easier variations, LLMs have not readily generalized this way. An LLM that could reliably solve $(9/5)x+32$ could not solve $(7/5)x+31$ (mccoy2023embersautoregressionunderstandinglarge). An LLM correctly defines an ABAB rhyming scheme but then generates poems violating its own definition (mancoridis2025potemkin). An LLM trained on “A is B” will not know “B is A” (berglund2023reversal) and LLMs struggle with logical puzzle variations (nezhurina2024alicewonderlandsimpletasks). Brittleness extends to presentation: LLMs answer multiple-choice questions correctly but fail when answer order is permuted (zong2024foolvisionandlanguage), succeed at programming with 0-based indexing but fail with 1-based indexing (wu2024reasoningrecitingexploringcapabilities), and show inconsistent performance on similar spatial reasoning tasks (MelanieMitchell(23)-Brittleness).
LLM errors align poorly with human intuitions precisely because we expect them to understand and misunderstand as humans do. Dellacqua(23) aptly characterize this as AI’s “jagged frontier” – a landscape where performance on similar tasks varies unpredictably. This jagged frontier captures why benchmark performance cannot be intuitively generalized to specific economic applications.
2.3.2 Do Large Language Models have World Models?
Perhaps this is too pessimistic, and this brittleness is superficial – noise around a deeper structural understanding embedded within LLMs. LLMs are trained on massive corpora containing rich information about reality, followed by extensive reinforcement learning. Perhaps LLMs have successfully learned “world models” – internal, structured representations of how the world works – that would enable reliable generalization despite occasional errors.
This hypothesis is an active research area focused on settings where researchers can both infer the LLM’s implicit world model and compare it to the truth (e.g., li2024emergentworldrepresentationsexploring; nanda2023emergentlinearrepresentationsworld; jylin04_2024_othellogpt; nikankin2025arithmeticalgorithmslanguagemodels). The evidence so far is at best mixed and at worst suggests LLMs may not learn generalizable world models. For example, vafa2024worldmodel trained a generative sequence model on turn-by-turn driving data from 12.6 million NYC taxi rides. While it predicts the next turn between locations with high accuracy, the authors reverse-engineered the model’s implicit map of Manhattan’s street grid: the implicit map bears no resemblance to the actual grid. High predictive accuracy did not require learning the underlying spatial structure. vafa2025has trained a generative sequence model on planetary orbits. Despite accurately predicting planetary movements, the model reveals no understanding of Newtonian physics. Rather than learning the gravitational law, it relies on task-specific heuristics that fail to generalize beyond its training distribution. Even advanced reasoning models exhibited similar failures.
Taken together, the evidence suggests that LLMs not only make errors, but these errors cannot be reliably predicted. They fail in ways misaligned with how humans generalize across tasks, nor can we assume these errors are noise around accurate world models. This creates our challenge. Since LLMs are impressive yet brittle tools, how should researchers incorporate their outputs into empirical research?
3 Prediction with Large Language Models
We begin with prediction problems – using text to predict economic outcomes. LLMs’ extensive training makes them natural candidates for this task, having already learned rich representations of language. Valid inference in prediction problems hinges on one condition: no leakage between the model’s training data and the researcher’s evaluation sample. While training leakage plagues benchmark evaluations in computer science, economists can manage it through careful model choice and research design.
@@ 3.1 The Researcher’s Prediction Problem
[…]
@@
where $N=\sum_{r}D_{r}$ is the number of text pieces collected. We would like to draw conclusions about the predictability of $Y_{r}$ from the text piece $r$ based on the LLM’s sample average loss. When is this valid?
Researchers face two sources of uncertainty about LLMs: What data was the LLM trained on? How does it generate output from text? We introduce two objects to formalize these uncertainties. The research context formalizes what we know (and do not know) about the LLM’s training data. The LLM’s guarantee summarizes high-level properties about how the model behaves without requiring exact knowledge of its internal workings. Together, these will allow us to derive interpretable conditions for using LLMs in prediction problems.
The research context $Q(\cdot)\in\mathcal{Q}$ is a joint distribution over the sampling indicators $(D,T)$. It encodes two distinct features. The sampling distribution over $D$ defines the researcher’s out-of-sample prediction problem – the collection of text pieces over which they want to evaluate the LLM’s predictive performance. The sampling distribution over $T$ captures the researcher’s uncertainty about the LLM’s training data. We make a technical assumption about the collection of research contexts $\mathcal{Q}$.
@@
Letting $t=(t_{\sigma_{1}},\ldots,t_{\sigma_{|\Sigma^{*}|}})^{\prime}$ denote the sampling indicators summarizing the LLM’s realized training dataset, all research contexts $Q(\cdot)\in\mathcal{Q}$ satisfy: (i) For all values $d$, $Q(D=d,T=t)=\Pi_{\sigma\in\Sigma^{*}}Q(D_{\sigma}=d_{\sigma},T_{\sigma}=t_{\sigma})$; and (ii) $\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}]=\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\mid T=t]$.
Assumption 1(i) states that sampling across strings is independent but not necessarily identically distributed. Assumption 1(ii) states that the researcher’s expected sample size does not depend on the LLM’s training corpus. Let $q^{T\mid D}_{\sigma}(t_{\sigma})=Q(T_{\sigma}=t_{\sigma}\mid D_{\sigma}=1)$ denote the conditional probability the string is sampled by the LLM’s training dataset given that it is sampled by the researcher. The marginal probabilities are $q^{T}_{\sigma}(t_{\sigma})=Q(T_{\sigma}=t_{\sigma})$ and $q^{D}_{\sigma}=Q(D_{\sigma}=1)$.
Our goal is to assess the LLM’s predictive performance over the population of text pieces defined by the research context $Q(\cdot)$:
$\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r;t))\right].$ (2)
This is summarizes whether the model’s predictive performance on the broader target population. As the number of economically relevant text pieces grows large (see Appendix C.1), we can characterize what the sample average loss converges to. Conditional on the LLM’s realized training dataset,
$\frac{1}{N}\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r;t))-\frac{1}{\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\mid T=t]}\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r;t))\mid T=t\right]\xrightarrow{p}0.$ (3)
Conditioning on the training dataset $t$ treats the LLM $\widehat{m}(\cdot;t)$ as a fixed mapping, avoiding strong assumptions about how the LLM was trained (e.g., how would it behave over counterfactual training datasets?). The sample average loss recovers our target quantity in Equation (2) if $\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r;t))\mid T=t\right]=\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r;t))\right].$
Recall the second source of uncertainty: researchers do not know precisely how the LLM generates outputs from text. Yet researchers often observe other high-level information about the LLM’s behavior, such as performance on benchmarks (e.g., ”achieves 88.7% on MMLU” or ”scores 1520 on the SAT”). We formalize this through a guarantee $\mathcal{M}$, a collection of possible text generators that captures what the researcher knows about the LLM’s behavior. The researcher only knows that $\widehat{m}(\cdot;t)\in\mathcal{M}$.
Given a guarantee $\mathcal{M}$ and a research context $Q(\cdot)$, we define when this workflow is valid for prediction problems.
Definition 1 (Prediction problem).